OpenStudy (anonymous):

TRIGNOMETRY If A and B are acute positive angles satisfying the equation 3 sin2 A + 2 sin2B = 1 and 3 sin 2A – 2 sin2B = 0, then find the value of (A + 2B).

OpenStudy (welshfella):

are all these double angles? 2A ad 2B?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

IN THE FIRST EQUATION THIS IS SQUARE'S NOT DOUBLE ANGLES

OpenStudy (anonymous):

3 sin^2 A +2sin^2 B = 1

OpenStudy (welshfella):

right - i was'nt sure about that.

OpenStudy (anonymous):

I THINK YOU NEED THESE EQUATIONS AGAIN

OpenStudy (anonymous):

3 sin^2 (A) + 2 sin^2(B) = 1 and 3 sin 2A – 2 sin2B = 0,

OpenStudy (anonymous):

I THINK THIS TIME IT'S MORE CLEAR

OpenStudy (anonymous):

IS IT

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

PLEASE SOLve and give me the whole,understand able solution

OpenStudy (anonymous):

i give you medal;

OpenStudy (welshfella):

i would help you but at the moment I can't see a way to do this.

OpenStudy (welshfella):

No sorry I cant think of a way to do this one.

OpenStudy (welshfella):

I'LL TRY IT OM WOLFRAMAPLHA

OpenStudy (welshfella):

http://www.wolframalpha.com/input/?i=3+sin^2+A+%2B+2+sin^2+B+%3D+1%2C+1+%2B+3+sin+2A+-+2+sin+2B+%3D+0+Solve+the+system That is Complex!!!! I think there might be an error in your question

OpenStudy (welshfella):

though I tried entering the criteria that angles nust be acute and positive and 'find A + 2B' but it ignored it.

OpenStudy (sparrow2):